OpenStudy (anonymous):

solve, g(p) as a comp of 3

OpenStudy (anonymous):

i guess it is f(x) x g(x) , f(x) = x^2 g(x) = 1/(x^2+4)

OpenStudy (agent0smith):

You could let one function be f(x) = x^2, and the other g(x) = x/(x+4) Then g(f(x)) = F(x), since if you plug in x^2 into g(x), you get F(x) There is other ways you can do it, too

OpenStudy (agent0smith):

What I said is correct. The other poster's was not a composition of functions.